OpenStudy (anonymous):

Find the equation of a circle in standard form where C(6, -2) and D(-4, 4) are endpoints of a diameter.

OpenStudy (anonymous):

first find the center, which is the middle of the two points. it is \[(\frac{6-4}{2},\frac{-2+4}{2})\] or \[(-1,1)\]

OpenStudy (anonymous):

dam i made a mistake. the center is not \[(-1,1)\] it is \[(1,1)\]

OpenStudy (anonymous):

bad arithmetic. now you need the radius from the distance formula \[r^2=(6-1)^2+(-2-1)^2=5^2+3^2=25-9=34\]

OpenStudy (anonymous):

so your equation is \[(x-1)^2+(y-1)^2=34\]